TCCON > Sites > JPL > Instrument History from August 2007 to June 2008

Instrument History from August 2007 to June 2008

The instrument currently at Lamont spent just under a year at JPL to measure the atmosphere simultaneously with OCO during its thermal vacuum testing.

June 2008

  • Pictures of the move out of JPL here
  • June 26 - Last day of measurements before packing up to move to Lamont
  • June 17 - Measured LASA and LASB and detector levels prior to move
  LASA LASB
  PK-PK   Zero PK-PK   Zero
Front -9.000V -1.063V -9.188V -7.250V 3.00V -9.188V
Back -8.000V -625mV -9.125V -6.250V 2.250V -9.250V

DC Values

InGaAs  : APT & AP2 = 1000; -0.700 VDC. When AP2=1300, -0.807VDC.
            : DC Counts: Amplitude: 4096; Position 149324 (1.3 mm field stop; high pass filter on)
            : DC Counts: Amplitude: 9139; Position 149324 (1.3 mm field stop; high pass filter open)
Si         : APT & AP2 = 1000; -0.163 VDC
            : APT = 1000; AP2 = 1300; -0.207 VDC
            : APT = 1300; AP2 = 1300; -0.366 VDC
            : DC Counts: Amplitude: -1106; Position 149326 (1.3 mm field stop; high pass filter on)
            : DC Counts: Amplitude: 264; Position 149326 (1.3 mm field stop; high pass filter open)

  • June 2 - Hercules motherboard died over the weekend (May 30-June1), and was replaced
    • Swapped InGaAs detectors

May 2008

  • May 28 - Switched InGaAs detector at 10:30 am
  • May 16 - Zeno: 969.9 hPa; Baro: 732.05 mmHg @ 23.2C

April 2008

  • April 30 - Zeno: 967.1 hPa; Baro: 730.03 mmHg @ 25.1C
  • April 9 - Zeno: 967.3 hPa; Baro: 730.5 mmHg @ 26C

March 2008

  • March 19 - Zeno: 964.5 hPa; Baro: 728.35 @ 18.9C

February 2008

  • February 25 - Zeno: 975.2 hPa; Baro: 735.5 mmHg @ 22C
    • OCO heliostat data to be recorded tomorrow
  • February 21 - Zeno: 970.5 hPa; Baro: 732.1 mmHg @ 18C
  • February 12 - Zeno: 973.3 hPa; Baro: 734.4 mmHg @ 24.4C
  • February 7 - Zeno: 973.8 hPa; Baro: 734.85 mmHg @ 24C
  • February 1 - Zeno: 972.1 hPa; Baro: 733.45 mmHg @ 19.8C (possible reading error)

January 2008

  • January 25 - Zeno: 974.9 hPa; Baro: 735.05 mmHg @ 17.5C
  • January 17 - Zeno: 970.6 hPa; Baro: 732.2 mmHg @ 21.3C
  • January 14 - Zeno: 978.6 hPa; Baro: 738.05 @ 21.2C
  • January 7 - Zeno: 968.7 hPa; Baro: 730.6 mmHg @ 20C
    • OCO heliostat measurements today

December 2007

  • December 19 - Zeno: 976.7 hPa; Baro: 736.3 mmHg @ 19C
  • December 3 - Zeno: 982.9 hPa; Baro: 740.3 mmHg @ 13.5C
    • Scanner error in the morning - IFS is cold (8C) - turned on heater
    • Restarted IFSloop

November 2007

  • November 30 - IFS turned off in preparation for power cut over the weekend
    • turned off IFSloop
    • shut down hercules
    • disconnected batteries
    • turned off lower, then upper power switch on IFS
    • durned off IFS power bar
    • turned off dome
    • suntracker - F10 x2
      • turned off control box and computer
    • turned off power bars, line stabilizers
    • UPS off
    • disconnected IFS
  • November 28 - removed InSb detector
    • Zeno: 974.9 hPa; Baro: 735.35 mmHg @ 22C
    • Note on door: Power cut planned for Saturday, Dec. 1, 2007 from 7am-12pm
  • October 20 - November 5: wildfires evident in spectra
  • November 6 - Zeno: 970.4 hPa; Baro: 731.35 mmHg @ 20C

October 2007

  •  InSb scans recorded Oct 25-26 (before alignment); Oct 27, 30 (after alignment)
  • October 26 - Zeno: 967.5 hPa; Baro: 729.85 mmHg @ 22C
  • October 25 - filled InSb at 1:29 pm
  • October 24 - filled InSb at 6:30 am, but TMA not reading
  • October 23-26 - major forest fires nearby
  • October 17 - Zeno: 968 hPa; Baro: 732.45 mmHg @ 23.5C
  • October 10 - Zeno: 972.0 hPa; Baro: 733.05 mmHg @ 21C
  • October 5 - Zeno: 965.6 hPa; Baro: 728.85 mmHg @ 22.8C

September 2007

  • September 27 - Zeno: 969.4 hPa; Baro: 731.2 mmHg @ 23.2C
  • September 24 - TVAC1 is finished.
  • September 17 - Zeno: 968.7 hPa; Baro: 730.85 mmHg @ 23.8C
  • September 5 - Zeno: 961.1 hPa; Baro: 721.75 @ 23C - corrected to 962.254 hPa

August 2007

  • August 30 - realigned suntracker on diode
  • August 23 - adjusted suntracker on diode
  • August 20 - taped index cards over opening in cylinder on roof to prevent direct sunlight from entering the CaF2 window directly
  • August 14 - dome retrained (but a cloudy day)
  • August 13 - recentered solar beam on entrance pinhole
  • August 9 - put paper constricting input to IFS between 10:35am and 11:51am
  • August 3 - dome facing NE, restarted IFSloop

July 2007

  • July 31 - Replaced tygon tubing on zeno sensor
    • Zeno: 967.2 hPa; Baro: 729.75 mmHg @ 25.5C - corrected to 967.92 hPa
  • Pictures of the move from Caltech to JPL here.
Tag page